Find the co-ordinates of the points on the curve Y=X^3-2x^2+3x-5 where its gradient is 2
1
Expert's answer
2019-09-19T10:37:30-0400

"\frac{dy}{dx}=3x^2-4x+3."

"\frac{dy}{dx}=2, so\; 3x^2-4x+3=2 \; or\; 3x^2-4x+1=0."

Thus, "x=\frac{1}{3}\;or\;x=1."

"y(\frac{1}{3})=-\frac{113}{27}, y(1)=-3."

Points on the curve where its gradient is 2: "(\frac{1}{3},-\frac{113}{27}) \; and\; (1,-3)."